# Defining A New Cross Reality: Digital Twins and Mixed Reality Worlds

### v1

@@ Abstract
Technological advances are redefining the relationship between physical and virtual space. Traditionally, when users engage in virtual reality (VR), they are completely cut off from the physical space; similarly, they are unable to access virtual experiences while engaged in physical activities. However, modern multi-platform metaverse environments allow simultaneous participation through mobile devices, creating new opportunities for integrated experiences. This study introduces the concept of "cross-reality lifestyles" to examine how users actively combine their physical and virtual activities. We identify three patterns of integration: 1) amplification: one space enhances experiences in the other; 2) complementary: spaces offer different but equally valuable alternatives; and 3) emergence: simultaneous engagement creates entirely new experiences. By analyzing commercial platforms, we create a technical framework that addresses content design, platform infrastructure, and device interfaces. This framework guides the development of cross-reality applications while demonstrating how metaverse technologies blur the traditional boundaries between physical and virtual experiences.
@@
The relationship between physical and virtual space is being redefined by advances in information technology. Since the era of video games and early social virtual worlds (such as Second Life), real and virtual space have been treated as mutually exclusive: users in virtual spaces were cut off from the information of physical space, while virtual experiences were disconnected from everyday life. In recent years, advancements in immersive virtual reality technology, offering a sense of presence and enabling embodied avatar experiences, have further enhanced the ways users interact with virtual environments [1]. VR creates a sense of being in a completely different place, enabling the simulation of extraordinary, non-everyday experiences. This separation shaped the technological development in two different directions. Digital twin technology [2] positioned virtual spaces as tools for simulating and optimizing real environments, establishing physical space as primary. In contrast, metaverse platforms [3] such as VRChat [^1] https://hello.vrchat.com/ and Cluster [^2] https://cluster.mu/en created virtual worlds independent of physical constraints, creating new venues for self-expression free [4]. These platforms prioritize virtual experiences while minimizing physical connections.
However, current trends suggest a shift beyond this dualistic division between real and virtual spaces. For example, younger generations routinely combine physical activities with virtual social connections, such as doing chores while staying connected through video calls [5]. Co-existence of activities in real and virtual spaces has expanded through metaverse development, especially as platforms adopt cross-device accessibility from mobile devices, such as VR-HMD and smartphones, and IoT device integration. For example, Roblox [^3] https://www.roblox.com/, one of the world’s largest non-immersive metaverse platforms added VR-HMD support in 2023, while VRChat, which is played by many VRHMD users,introduced smartphone access in 2024. As activities in virtual spaces have grown in popularity and diverse access routes to these spaces have been developed, users who simultaneously engage in both real and virtual spaces have emerged. As seen in Fig. 1, users can participate in metaverse events while doing something in real spaces [^4] Cluster Music Jam Vol.1. https://cluster.mu/e/96cf9bff-5cd5-406b-ac50-dc51e9ef08a0. While previous studies have primarily focused on the master-slave relationship that exists between the two spaces, such as real-world VR training [6] or user activity in metaverse platforms [4], these emerging patterns of simultaneous engagement remain understudied.
@@
In this paper, we introduce the concept of “cross-reality life experiences” to describe emerging lifestyle patterns that arise from the seamless fusion of physical space and the metaverse. We define these experiences as novel forms of activity where neither space takes precedence over the other, and life in both realms integrates into a unified activity. These experiences encompass everything from enhancing real-world capabilities through VR to crafting personalized entertainment by harmonizing the interplay between physical and virtual actions.
We categorize these cross-reality life experiences into three integration patterns: 1) amplification, 2) complementary, and 3) emergence. Using examples from virtual platforms and communities, we analyze their distinct characteristics. Additionally, by mapping representative commercial metaverse platforms to technical requirements and integration patterns, we propose a framework for designing effective cross-reality experiences. This research offers practical insights for developers and designers aiming to create impactful and meaningful cross-reality applications.
@@
Milgram et al.’s Reality-Virtuality Continuum [7] positioned real and virtual spaces as poles on a spectrum, with Augmented Reality (AR), which augments physical space with virtual information, and Augmented Virtuality (AV), which augments virtual space with physical elements, filling the space between these poles. While this framework focused primarily on perceptual experiences, our research extends it to everyday activities. This extension emphasizes active user agency in combining physical and metaverse spaces.Users dynamically adjust their level of engagement based on situation and intent, moving fluidly between spaces. This concept of "cross-reality life experience" extends Mixed Reality into everyday life.
Figure 2: Patterns of integration between physical reality (R) and metaverse (M) spaces. The diagram illustrates three distinct patterns: (1) Amplification, where one space catalyzes value enhancement in the other through both reality-first and metaverse-first approaches; (2) Complementary, where M provides alternative options to overcome R’s inherent limitations while maintaining R’s core strengths; and (3) Emergence, where parallel engagement in both spaces creates qualitatively new experiences beyond simple combination.
We analyze these emerging lifestyle patterns through two dimensions: "temporal relationship," which examines whether one space takes precedence or if they coexist simultaneously, and "value interaction pattern," which explores how activities across spaces generate new forms of value. Based on these dimensions, we identify three types of integration: amplification, complementary, and emergence (Figure 2). The following sections examine each type with relevant literature.
Amplification-type Integration
Amplification-type integration involves enhancing the experiential value of either metaverse (M) or real space (R) through the intervention of the other. This pattern follows a clear temporal sequence, manifesting as either Reality-First or Metaverse-First Amplification.
@@
Reality-first amplification builds upon experiences in real space, where metaverse elements enhance the inherent value. This pattern is characterized by how partial participation in the metaverse serves as a catalyst to qualitatively enhance real-world experiences. This reality-first amplification has traditionally been considered in the context of digital twins and VR training, and Slater and Sanchez-Vives provide a comprehensive overview [6].
This pattern has shown significant effects across multiple domains. Medical training, particularly surgical education, uses virtual spaces to develop complex skills in reality while ensuring safety [8, 9, 10]. VR/AR applications have also been extensively studied in tourism [11]. Importantly, virtual travel experiences complement rather than replace physical travel, increasing motivation with VR travel or enhancing real visits with AR.
In physical activities such as rehabilitation and sports, VR enables graded skill development through calibrated challenges [12]. Virtual environments slightly above current skill levels promote successful experiences that enhance real-world performance [13]. These virtual achievements build self-efficacy, which translates into improved physical outcomes.
Researchers in psychology and neuroscience has revealed new insights into real-body ownership and perception by using VR as a tool [14, 15, 16]. Research also demonstrates that embodying avatars of different races can reduce implicit bias and develop empathy [17, 18].
@@
Metaverse-first amplification occurs when physical activities deepen the value originally created in virtual spaces. Classic examples include "anime pilgrimages," where fans visit real-world locations featured in anime or games to deepen their connection to the fictional narratives [19]. Similarly, offline "meet-ups" of virtual communities, such as online gaming groups and social networking sites, strengthen virtual bonds through physical interaction [20].
More contemporary examples include fan activities surrounding virtual YouTubers (VTubers) [21], where physical events and merchandise enhance virtual engagement. Metaverse platforms also leverage this pattern through events, connecting user-generated content (UGC) in virtual space with physical influence via streamer commentary [^5] Cluster Game Revolution Eve, 2024. https://geekaku2024.cluster.mu/, fostering broader metaverse community engagement and creative participation.
Complementary-type Integration
Complementary integration restores experiential wholeness by combining physical and metaverse options. This pattern features "simultaneously selectable states" where both spaces exist as equal alternatives, allowing users to choose optimal combinations for their needs.
Complement via Teleexistence
Telepresence robots in customer service exemplify this integration by reducing spatial and temporal constraints [22]. Virtual platforms, such as teleconferencing tools, complement face-to-face interactions by providing alternative modes of engagement. Additionlay, "silent study sessions" on Zoom provide ambient presence without social pressure [23], while screen sharing capabilities can enhance productivity beyond traditional meetings.
The COVID-19 pandemic has accelerated the adoption of virtual platforms such as Gather and Spatial Chat [24] for education and conferencing. These lightweight, browser-based 2D virtual environments, even without complex 3D capabilities, effectively complement physical spaces with unified spatial experiences.
Complement via Avatars
@@
On the other hand, Hatada et al. introduced a avatar system in a cafe where individuals with disabilities provide customer service using telepresence robots [28]. The robots’ standardized appearance made it difficult for employees to express their individuality, so the system incorporated customizable virtual avatars as an additional option for customer interactions. This enabled employees to seamlessly switch between physical telepresence robots and virtual avatars, allowing them to provide service from both the real and virtual spaces. The system empowered employees to engage more dynamically and authentically, creating a richer and more fulfilling customer service experience.
Emergence-type Integration
Emergent integration occurs when activities in the metaverse and physical spaces take place simultaneously and interact. Unlike complementary integration, which provides various options to users by switching across the metaverse and physical spaces, emergent integration involves the persistent coexistence and active combination of both spaces. This simultaneous integration creates a unified, multifaceted experience that would be impossible to achieve in either space alone.
Two technological advances enable this integration: the maturation of social VR communities and the widespread availability of hands-free VR devices, including VR-HMDs and smartphones. These developments make simultaneous engagement in both spaces technically feasible for general users.
Emergence with Incomplete Social Presence
The HCI and CSCW community has identified "incomplete" social presence as a significant pattern in which users modulate their level of digital engagement. Research shows that teenagers actively value this partial social presence, such as multitasking while participating video calls [5]. "Study with me" videos on YouTube Live demonstrate how users create personalized digital social environments that balance engagement with controllable peer pressure [29].
@@
Future of Emergence-type Integration
Advances in cognitive psychology with VR have expanded our understanding of embodiment and broadened the possibilities for emergent integration. AR avatar [31], which overlays virtual avatars on physical bodies through AR displays, brings the Proteus Effect [32], the observation that avatar characteristics influence users’ behavior and self-perception, into physical reality. This integration creates qualitative changes in cognitive and behavioral patterns that go beyond mere visual changes.
Research into the control of multiple virtual bodies [33] suggests possibilities beyond simple metaverse-reality combinations (M and R), introducing simultaneous engagement in multiple virtual spaces while maintaining physical presence (M1, M2, … and R).
TECHNICAL FRAMEWORK FOR CROSS-REALITY LIFESTYLE
This section presents an analytical framework for implementing cross-reality integration in metaverse environments. We first examine fundamental technical requirements across key dimensions, then analyze current platform implementations, and explore how UGC and multi-platform capabilities enable emergence-type integration.
Technical Requirements and Architecture
We establish three fundamental dimensions that characterize metaverse environments and organize their technical requirements:
  • Platform: Core technical infrastructure including multi-device support, real-time communication systems, scalability for multiple users, unified 2D/3D space management, and persistent data storage.
  • Device & Interface: Technologies bridging physical and virtual spaces through sensing, motion capture, VR/AR devices, and physical feedback systems.
Table 1 systematically organizes the relationship between these dimensions and the integration types by categorizing the technical requirement from examples shown in the previous section. For instance, while amplification-type integration requires high-precision training simulations for specific physical tasks, emergence-type integration demands functionalities supporting loose synchronous interaction between both spaces such as automated avatar control.
Table 1: Technical Requirements Matrix for Cross-Reality Integration: Mapping of Technical Dimensions against Integration Patterns.
Dimension Amplification Complementary Emergence
Contents Reality-first: [ ] High-precision training simulations for specific physical tasks (e.g., medical procedures, sports) Structured learning content that maps directly to real-world actions Context-aware guidance systems that enhance physical performance Metaverse-first: [ ] Virtual community or event (e.g., VTuber concerts) UGC management and promotion Teleconference: [ ] Meetings, seminars with stable UIs (e.g., screen sharing, easy-to-use controls) Tele-existence: [ ] Remote robot interaction or field monitoring with interactive control Avatar: [ ] Flexible avatar representation Privacy-conscious social interaction tools (e.g., Voice changer) [ ] A system without excessive manipulation in VR (e.g. automation of avatar movements and reactions) Flexible UGC tools enabling spontaneous cross-reality activities VR icon indicating physical status
Platform Reality-first: [ ] High-performance physics simulation engines Precise spatial tracking and mapping Metaverse-first: [ ] Scalable content distribution systems Cross-platform compatibility (VR, PC, mobile) Robust UGC management [ ] Low-latency, stable connections Reliable multi-user session management Compatibility across 2D/2.5D/3D environments via universal API Scaling for varying group sizes Protocols for device integration [ ] Cross-platform compatibility Always-on infrastructure supporting partial presence Dynamic resource allocation based on participation status Context-aware state synchronization
Device &
Interface
Reality-first: [ ] High-precision tracking devices Specialized input/output hardware and its integration (sensors and actuators) Metaverse-first: [ ] Consumer-grade VR/AR devices Accessible mobile interfaces Teleconference: [ ] Standard video/audio equipment Tele-existence: [ ] Remote operation control devices Environmental sensing Avatar: [ ] Motion capture for expressive virtual avatars (VTuber or streamer use) Head/face tracking or lip-sync for emotion expression [ ] Lightweight, comfortable devices for extended use Switching between Reality and Virtuality (e.g. video see-through) Multi-avatar or multi-world participation (e.g. controllers for switching input among instances) Integrated monitoring systems that track physical condition
@@
These platforms focus on workplace simulation and digital twin creation for the manufacturing and construction industries. They integrate CAD data and real-time sensor information to optimize plant layouts and verify operational processes. Technical requirements include precise industrial data compatibility, secure high-performance infrastructure, and precise control environments through AR/VR devices. Content priorities include accurate physical simulation with real-time capabilities, while the platform must provide enterprise-class reliability and scalability. These platforms primarily support reality-first amplification through sensor and actuator integration, and complementary integration through remote robot operation.
@@
Originally developed as gaming platforms with predefined goals, these systems have evolved towards metaverse functionality through enhanced UGC capabilities and large-scale events. Technical requirements include advanced 3D graphics processing, support for large concurrent connections, and optimization across multiple devices (PC, game consoles, mobile). Many platforms offer proprietary development environments and marketplaces for UGC. In terms of cross-reality integration, these platforms demonstrate strong metaverse-first amplification through large-scale virtual events. Social interactions within games increasingly extend to physical activities (offline meetings, merchandise sales). While emergent community development remains limited, increasing UGC capabilities may lead to spontaneous use cases.
@@
These platforms focus on communication and world editing through 3D avatars, supporting diverse applications such as music performances, academic conferences, and community events. Technical requirements include high quality 3D avatar rendering, multi-platform support (VR/PC/mobile), and advanced body tracking technologies such as motion capture. They offer significant UGC freedom and encourage active user creation of avatars and worlds. These platforms demonstrate the diverse cross-reality integration patterns, including metaverse-first amplification through VR events extending to physical activities, complementary integration through avatar-based customer service and meetings, and emergent integration exemplified by unique cultural phenomena like VR sleep.
UGC and Multi-platform Development
Analysis shows that UGC freedom and multi-platform support play a critical role in deepening cross-reality experiences. Platforms with greater UGC freedom enable greater user creativity, leading to unexpected emergent activity. Multi-device support reduces the barriers to participation, thereby facilitating the formation of emergent communities. However, mobile platforms often face UGC limitations due to device constraints, presenting a challenge for cross-reality experience development. Future development will need to maintain experience consistency across devices, while allowing for graded UGC deployment that leverages the characteristics of each platform.

### v2

@@ Abstract
Technological advances are redefining the relationship between physical and virtual space. Traditionally, when users engage in virtual reality (VR), they are completely cut off from the physical space. Similarly, they are unable to access virtual experiences while engaged in physical activities. However, modern multi-platform metaverse environments allow simultaneous participation through mobile devices, creating new opportunities for integrated experiences. This study introduces the concept of "cross-reality lifestyles" to examine how users actively combine their physical and virtual activities. We identify three patterns of integration: 1) Amplification: one space enhances experiences in the other; 2) Complementary: spaces offer different but equally valuable alternatives; and 3) Emergence: simultaneous engagement creates entirely new experiences. We propose the ACE Cube framework that analyzes these patterns as continuous characteristics, and by integrating this analysis with technical requirements of commercial platforms, we provide practical guidelines for platform selection, technical investment prioritization, and cross-reality application development.
@@
The relationship between physical and virtual spaces is being redefined by advances in information technology. Since the era of video games and early social virtual worlds like Second Life, these spaces have been treated as mutually exclusive: users in virtual spaces were isolated from the physical information, while virtual experiences remained disconnected from everyday life. Recent developments in immersive VR technology have intensified this separation by creating a stronger presence and an embodied avatar experience [1]. VR creates a sense of being in a completely different place, enabling the simulation of extraordinary, non-everyday experiences. This separation shaped the technological development in two different directions. Digital twin technology [2] positions virtual spaces as tools for simulating and optimizing real environments, establishing physical space as primary. In contrast, metaverse platforms [3] such as VRChat [^1] https://hello.vrchat.com/ and Cluster [^2] https://cluster.mu/en prioritize virtual worlds independent of physical constraints, creating new venues for unconstrained self-expression [4].
However, current trends suggest a shift beyond this dualistic division. For example, younger generations routinely blend physical activities with virtual social connections, such as doing chores while staying connected through video calls [5]. The coexistence of activities in real and virtual spaces has expanded through metaverse development, especially as platforms adopt cross-device accessibility from mobile devices, such as VR-HMD and smartphones, and IoT device integration. For example, Roblox [^3] https://www.roblox.com/, one of the world’s largest non-immersive metaverse platforms added VR-HMD support in 2023, while VRChat, which is played by many VR-HMD users, introduced smartphone access in 2024. As diverse access routes to these spaces have been developed, users simultaneously engage in both spaces, participating in metaverse events while performing physical activities (Fig. 1) [^4] Cluster Music Jam Vol.1. https://cluster.mu/e/96cf9bff-5cd5-406b-ac50-dc51e9ef08a0. Unlike previous studies that have focused mainly on the master-slave relationship that exists between the two spaces, such as real-world VR training [6] or user activity on metaverse platforms [4], these emerging patterns of simultaneous engagement remain understudied.
@@
In this paper, we introduce the concept of “cross-reality life experiences” to describe emerging lifestyle patterns from seamless physical-metaverse fusion. We define these experiences as novel activities where neither space dominates and life in both realms integrates into unified activity. To overcome traditional virtual-real opposition, we reinterpret these as interaction patterns between multiple "verses"—different implementations of the same activity space.
We analyze these cross-reality life experiences as integration patterns between different types of verses (e.g., physical and metaverse) through three dimensions: Amplification, Complementarity, and Emergence. Rather than considering them as discrete categories, we conceptualize these dimensions as continuous characteristics, each of which is exhibited by experiences that vary in intensity across all dimensions. We map these experiences within a three-dimensional ACE cube and position commercial metaverse platforms alongside their technical requirements. This establishes connections between the experiences’ integration patterns and technical implementation. This multidimensional framework transcends the physical-virtual binary opposition and provides practical guidelines for selecting platforms, prioritizing technical investments, and developing cross-reality applications.
Figure 2: Patterns of integration between verse 1 (V1, physical reality) and verse 2 (V2, metaverse) spaces. The diagram illustrates three distinct patterns: (1) Amplification, where one space catalyzes value enhancement in the other through both reality-first and metaverse-first approaches; (2) Complementary, where the one verce provides alternative options to overcome anoter verse’s inherent limitations while maintaining another verse’s core strengths; and (3) Emergence, where parallel engagement in both spaces creates qualitatively new experiences beyond simple combination.
UNIFIED ANALYSIS THROUGH ACE CUBE
Milgram et al.’s Reality-Virtuality Continuum [7] positioned physical and virtual spaces as opposing poles on a one-dimensional spectrum, with Augmented Reality (AR) augmenting physical space with virtual information, and Augmented Virtuality (AV) augmenting virtual space with physical elements filling the intermediate space. While this framework focused primarily on perceptual experiences, our research extends it to everyday activities.
Zeltzer’s AIP cube [8] pioneered three-dimensional VR system analysis. The AIP cube is a three-dimensional framework that evaluates technical characteristics of VR systems along three axes: Autonomy, Interaction, and Presence axes, with continuous values from 0 to 1. The coordinate (1, 1, 1) represents the "ultimate VR system".
Figure 3: The ACE Cube Framework is used to analyze cross-reality experiences. This three-dimensional framework maps cross-reality experiences along the Amplification (A), Complementarity (C), and Emergence (E) axes.
We propose a new three-dimensional framework called the "ACE cube" that analyzes value creation in users’ daily-life experiences, while the AIP cube measures technical system capabilities. We analyze integration patterns between physical and metaverse spaces as continuous characteristics along three dimensions (Fig. 2):
  • A-axis (Amplification) measures how experiences in one space enhance capabilities in another. For example, VR training improves real surgical skills, and real-life events deepen fan activities for virtual YouTubers (VTubers). From a temporal perspective, amplification-type integration involves a clear sequence in which activities in one space enhance activities in another.
  • C-axis (Complementary) quantifies how one space complements the constraints and limitations of the other. This restores experiential wholeness through metaverse-physical fusion, which was previously limited by single-space constraints due to the absence of metaverse options. From a temporal perspective, complementary-type integration is characterized by "simultaneously selectable" states. Multiple spaces exist as equal alternatives, enabling the selection of optimal combinations according to the situation.
  • E-axis (Emergence) measures qualitatively new values that arise from simultaneous multi-space experiences that exceed simple summation. From a temporal perspective, emergent-type fusion features "persistent simultaneity." While complementary integration recovers wholeness through switching spaces, emergent integration creates new experiences through parallel engagement.
Each cross-reality experience exhibits specific intensities along these axes, represented as continuous coordinates within the ACE cube (Fig. 3). We map experience cases and position existing platforms within this cube. The coordinate (1, 1, 1) represents the "ultimate metaverse platform" that supports all types of experiences.
Spatial Relationship Reconstruction Through the Verse Concept
Traditional analyses have assumed qualitative differences between "real" and "virtual" spaces. However, recent developments in the metaverse reveal that, although technical requirements differ, experiences across spaces are increasingly similar.
To overcome the "virtual-real" opposition, we reinterpret it as an interaction pattern between multiple "verses." Physical and metaverse spaces are different implementations of the same "verse," or activity space. For example, considering hosting a VTuber event, the physical implementation requires securing a venue, setting up a stage, and installing lighting and sound. In contrast, the metaverse implementation requires designing a virtual venue and constructing a 3D stage with virtual lighting and sound. Although the technologies differ, the essential structure of constructing event space remains identical.
From this perspective, there are multiple "verses" with different implementation technologies that exist in parallel. Each verse has specific constraints and affordances: physical verses face gravity and physical laws but provide tactile feedback and direct embodiment. In contrast, metaverses transcend physical constraints, offering expressive freedom and global accessibility.
The "verse" concept provides a neutral way to describe traditional relationships as "interverse interaction patterns." For example, amplification experiences essentially become uniform value enhancements through interverse interactions. Therefore, we represent both physical and metaverse activities as the same "interverse amplification" phenomena, whether physical activities improve metaverse capabilities or vice versa. Furthermore, this concept supports future multi-metaverse integration analysis. For example, when avatar creation skills in VRChat enhance event hosting capabilities in Cluster, this can be analyzed as "Metaverse $_{1}\rightarrow V$2" amplification.
@@
Figure 4: Cross-reality experiences and platform distribution in the ACE Cube. Two-dimensional projections show the A-C (left) and A-E (right) planes. Examples of experiences and commercial platforms are mapped onto these planes. When mapping the platforms, a yellow curve appeared on the A-C plane, indicating that existing platforms have adopted a clear strategic differentiation aligned with the A-C trade-off. However, social VR platforms demonstrate superior emergent capabilities on the A-E plane compared to platforms in other categories.
Figure 4 shows examples of cross-reality experiences mapped onto the ACE Cube. Three researchers created this ACE map by surveying over 200 official event cases on the Cluster platform and related work. Currently, this analytical method relies on the judgment of researchers. However, developing more objective mapping methods is a topic for future research. Based on this map, the following sections examine each type with relevant literature.
Amplification-oriented Experiences
Amplification-type integration involves enhancing the experiential value of one verse through the intervention of the other.
@@
Reality-first amplification builds on experiences in real space, where metaverse elements enhance the inherent value. This pattern is characterized by how partial participation in the metaverse serves as a catalyst to qualitatively enhance real-world experiences. This reality-first amplification has traditionally been considered in the context of digital twins and VR training, and Slater and Sanchez-Vives provide a comprehensive overview [6].
This pattern has shown significant effects across multiple domains. Medical training, particularly surgical education, uses virtual spaces to develop complex skills in reality while ensuring safety [9, 10, 11]. In contrast, functions that complement real surgical techniques are limited, and training and practice do not occur simultaneously. Thus, we plot them on high A and low C and E values.
VR/AR tourism applications combine travel experiences in a virtual environment with real tourism [12]. While these applications primarily serve amplification roles through AR-enhanced physical experiences and VR-improved travel motivation, they also serve complementary roles in overcoming the distance, cost, and time constraints of the physical world.
In psychology and neuroscience fields, VR research on body ownership has advanced, revealing flexibility in self-recognition that was difficult to observe with traditional methods [13, 14, 15]. Metaverse experiments deepen our understanding of the physical world. For example, researchers have reported that embodying avatars of different races can reduce prejudice and develop empathy in the physical world [16, 17].
AR avatar [18], which overlays virtual avatars on physical bodies through AR displays, brings the Proteus effect [19], the observation that avatar characteristics influence user behavior and self-perception, into the physical verse. The Proteus effect allows for the direct application of improved self-perception and behavioral patterns from the metaverse to the physical world. It also enables self-expression embodiment that would otherwise be impossible in physical environments through AR technology.
@@
Metaverse-first amplification involves strengthening the values inherent in metaverse by adding physical elements. The most basic example is anime pilgrimage, a cultural phenomenon that involves visits to real locations featured in anime and video games [20], which more strongly amplify the reality of metaverse experiences. In addition, offline "meet-ups" that deepen bonds formed in online game communities in the real world exemplify the metaverse space relationships being strengthened through physical interaction [21].
Contemporary developments of this amplification pattern include "oshi-katsu" (fan activities) supporting VTubers and streamers [22]. These activities have evolved into real-world behaviors such as participating in physical events and purchasing merchandise, which amplifies the fan experience.
Complementary-oriented Experiences
Complementary integration restores experiential wholeness by combining physical and metaverse options. Telepresence robots in customer service exemplify this integration by reducing spatial and temporal constraints [23]. Virtual platforms, such as teleconferencing tools, complement face-to-face interactions by providing alternative modes of engagement.
@@
Hatada et al. introduced an avatar system in a cafe where individuals with disabilities provide customer service using telepresence robots [27]. The robots’ standardized appearance made it difficult for employees to express their individuality, so the system incorporated customizable virtual avatars as an additional option for customer interactions. This enabled employees to seamlessly switch between physical telepresence robots and virtual avatars, allowing them to provide service from both the real and virtual spaces. The system empowered employees to engage more dynamically and authentically, creating a richer and more fulfilling customer service experience.
Alternative means of face-to-face interaction through virtual environments create usable options simultaneously for real and virtual, enabling optimal combinations according to the situation. "Silent study sessions" on Zoom provide ambient presence without social pressure [28], while screen sharing capabilities can enhance productivity beyond traditional meetings.
Emergence-oriented Experiences
Emergent integration occurs when activities in the multiple (e.g., real and virtual) verses take place simultaneously and interact. Two technological advances enable this integration: the maturation of social VR communities and the widespread availability of hands-free VR devices, including VR-HMDs and smartphones. These developments make simultaneous engagement in both spaces technically feasible for general users.
The HCI and CSCW community has identified "incomplete" social presence as a significant pattern in which users modulate their level of digital engagement. Research shows that teens actively value this partial social presence, such as multitasking while participating in video calls [5]. "Study with me" videos on YouTube Live demonstrate how users create personalized digital social environments that balance engagement with controllable peer pressure [29].
@@
The ultimate research on such emergent experiences includes multiverse body control [31]. This involves controlling bodies in multiple metaverses while maintaining a presence in the physical world. Having a single consciousness exist in multiple verses creates revolutionary experiences and suggests the potential transformation of traditional concepts of embodiment and spatiality.
@@
These platforms focus on workplace simulation and digital twin creation for the manufacturing and construction industries. They integrate CAD data and real-time sensor information to optimize plant layouts and verify operational processes. Technical requirements include precise compatibility with industrial data, secure high-performance infrastructure, and precise control environments through AR/VR devices. Content priorities include accurate physical simulation with real-time capabilities, while the platform must provide enterprise-class reliability and scalability. These platforms primarily support reality-first amplification through sensor and actuator integration, and complementary integration through remote robot operation.
@@
While these systems were originally developed as gaming platforms with predefined rules, they have evolved toward metaverse functionality through enhanced user-generated content (UGC) capabilities and large-scale events. Technical requirements include advanced 3D graphics processing, support for a high number of concurrent connections, and optimization across multiple devices, such as PCs, gaming consoles, and mobile devices. Many platforms offer their own development environments and marketplaces for UGC.
From a cross-reality integration perspective, these platforms specialize in complementary functions. For example, Minecraft and Roblox provide opportunities for large-scale construction and creation of structures that transcend physical laws and would be impossible in reality. These opportunities are difficult to achieve through traditional classroom instruction and serve as entry points for programming education. Additionally, the creative skills gained through gaming partially transfer to real-world capabilities; however, this remains moderate due to gaming’s primary focus on gaming experiences. This focus of the platform design also limits the creation of emergent experiences that occur simultaneously with physical activities.
@@
These platforms focus on communication and world editing through 3D avatars, supporting diverse applications such as music performances, academic conferences, and community events. Technical requirements include high-quality 3D avatar rendering, multi-platform support (VR/PC/mobile), and advanced body tracking technologies such as motion capture. They offer significant UGC freedom and encourage active user creation of avatars and worlds. Compared to other platforms, these platforms demonstrate superior capabilities in creating emergent experiences. These platforms host many high-E, or "while participating," experiences, including VR sleep phenomena and virtual music events combined with physical activities.
Figure 5: Technical requirements mapping in the ACE cube. Technical requirements are classified as content (circles), platform infrastructure (diamonds), or devices and interfaces (triangles) across projections A-C and A-E. The yellow lines in both planes are in the same position as in Fig. 4 They represent the technical requirements for realizing cross-reality experiences, as well as platform selection and mapping. In the A-C plane, the yellow line help select platform technology elements according to experience A-C trade-offs. In the A-E plane, the yellow line show that technologies positioned along the E-axis provide emergent experiences through supporting platforms.
TECHNICAL FRAMEWORK FOR CROSS-REALITY LIFESTYLE
To clarify the relationship between experience cases and technical requirements, we analyze the technical requirements for a cross-reality experience in the ACE cube.
These technical requirements are classified into three layers:
  • Platform Infrastructure: Core technical infrastructure including multi-device support, real-time communication systems, scalability for multiple users, unified 2D / 3D space management and persistent data storage.
  • Device & Interface: Technologies bridging physical and virtual spaces through sensing, motion capture, VR / AR devices and physical feedback systems.
Technical Requirements Mapping in ACE Cube
Figure 5 shows the ACE coordinate values for each technical requirement. This analysis clarifies the technical conditions necessary to realize specific experience domains.
  • High-A technologies (e.g., high-precision physics and spatial processing, high-precision tracking and sensing, and cross-platform compatibility) significantly enhance understanding and predictive capabilities through precise reality reproduction, supporting capability enhancement experiences such as medical training and VR rehabilitation.
  • High-C technologies (e.g., video communication, unified APIs and protocols, and security and privacy protection) serve as direct substitutes for physical face-to-face interaction by resolving distance, time, and technical compatibility constraints and enabling complementary experiences such as teleconferencing.
  • High-E technologies (e.g., avatar and motion automation, always-on infrastructure, mobile access, and context-aware synchronization systems) automate virtual presence during physical activities, enabling "parallel participation" and providing the basis for emergent experiences.
Platform Design Guidelines through Technology Integration
A comparison of the ACE analysis of existing platforms and the mapping of technical requirements reveals the factors that differentiate platforms from each other.
Social VR platforms significantly surpass others in their ability to support emergent experiences due to their integrated implementation of advanced technologies, such as avatar and motion automation, always-on infrastructure, and context-aware synchronization systems. These technologies provide the foundation for users to create new and unexpected experiences.
Industrial metaverse platforms specialize in high-precision simulation by focusing on advanced technologies, such as high-precision physics, spatial processing, and specialized input/output devices. These technologies enable experiences that significantly improve real-world operational efficiency. Teleconferencing tools focus on video communication technologies and security and privacy protection to serve as direct substitutes for physical constraints.
Emerging technologies, such as lightweight mixed reality (MR) devices, enable improved sustained use capabilities by reducing the burden of wearing them and by realizing new lifestyles that integrate virtual elements into daily life through extended, simultaneous experiences. These technologies are expected to pioneer new experience domains by integrating traditionally exclusive high-ACE regions.
This analysis helps developers identify the technical requirements for the target experience domains and formulate strategies to differentiate their platforms from existing ones. Developing platforms that approach the level of ACE Cube (1, 1, 1) requires integrated development in all technology domains.